OpenStudy (anonymous):

The circumference of a circular garden is 87.92 feet. What is the radius of the garden? Use 3.14 for (pie symbon inserted here) and do not round your answer.

OpenStudy (anonymous):

87.92=2*pi*r, plug in pi and solve for r

OpenStudy (anonymous):

Circumference C = 2*pi*r. To solve, let C = 87.92 and let pi = 3.14. Then you will be able to solve for r (the radius). Good luck!

OpenStudy (anonymous):

So, 87.92=2(3.14) 87.92/2 ?

OpenStudy (anonymous):

1. C = 2*pi*r 2. 87.92 = 2*(3.14)*r 3. 87.92/(2*3.14) = r 4. r = radius

OpenStudy (anonymous):

This makes no sence

OpenStudy (anonymous):

The equation for circumference is C = 2*pi*radius. You are given the circumference (C) and we know that pi = 3.14. The only variable you have left to find using simply math is the radius r. Does that help?